Apply the method of variations of parameters to solve the following differential equations: y" +a^2y=cosec ax
1
Expert's answer
2020-02-12T11:49:48-0500

"y''+a^2y=\cosec ax\\
y''+a^2y=\frac{1}{\sin ax}\\"

Solution can be found in the form

"y=y_1+y_2"

1) "y_1" is solution of equation

"y''+a^2y=0\\
\lambda^2+a^2=0\\
\lambda_1=ia, \lambda_2=-ia,\\
f_1=\cos ax, f_2=\sin ax\\"

then

"y_1=c_1 \cos ax+c_2 \sin ax"

2) solution y can be found by the method of variation of constants in the form

"y=c_1(x)\cos ax+c_2(x) \sin ax"

functions "c_1(x)" and "c_2(x)" can be found from the system of equations

"\left\{
\begin{matrix}
 c'_1(x) \cos ax+c'_2(x) \sin ax =0\\
 c'_1(x)(- a\sin ax)+c'_2(x) a \cos ax =\frac{1}{\sin ax}
\end{matrix}
\right."


"\Delta=\left|
\begin{matrix}
 \cos ax & \sin ax\\
 -a \sin ax & a\cos ax
\end{matrix}
\right|=\\
=a\cos^2 ax+a \sin^2 ax=a\\
\Delta_1=\left|
\begin{matrix}
 0 & \sin ax\\
 \frac{1}{\sin ax} & a\cos ax
\end{matrix}
\right|=-1\\
\Delta_2=\left|
\begin{matrix}
 \cos ax &0\\
 -a \sin ax & \frac{1}{\sin ax}
\end{matrix}
\right|= \frac{\cos ax}{\sin ax}\\
c'_1(x)=\frac{\Delta_1}{\Delta}=\frac{-1}{a}, \\
c'_2(x)=\frac{\Delta_2}{\Delta}=\frac{1}{a} \frac{\cos ax}{\sin ax}, \\
c_1(x)=\int \frac{-1}{a} dx=\frac{-x}{a} + k_1\\
c_2(x)=\int\frac{1}{a} \frac{\cos ax}{\sin ax}dx=\\
=\frac{1}{a^2} \int\frac{d(\sin ax)}{\sin ax}=\frac{1}{a^2}\ln|\sin ax|+k_2\\
k_1, k_2 \, are\,constants"

Then

"y=\left(\frac{-x}{a} + k_1\right)\cos ax + \left(\frac{1}{a^2}\ln|\sin ax|+k_2\right)\sin ax"